OpenStudy (anonymous):

Kenya is reserving a room in a hotel in France where they use euros (€). The room charge is €75. Suppose the conversion rate is €1 = $1.298. What is the cost of the room in dollars?

OpenStudy (jdoe0001):

what do you think?

OpenStudy (anonymous):

i don't understand how to answer the question ..

OpenStudy (anonymous):

i don't understand the concept basically

OpenStudy (jdoe0001):

say you go to a hotel, and you have room is €75 now you only have US dollars in your wallet so you go to the local exchange store to change your dollars for €75 how many dollars would you have to give to get €75 ?

OpenStudy (anonymous):

58 dollars?

OpenStudy (jdoe0001):

or lemme put it this way if the exchange rate between is €1 = $1.298 how many dollars do you have to give for say €3 ?

OpenStudy (anonymous):

i dont kno .... im struggling here ...

OpenStudy (anonymous):

plz help

OpenStudy (jdoe0001):

do you know what -> €1 = $1.298 <- mean?

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

€1 = $1.298 <--- means, for every Euro you'd want, you have to give 1.298 US dollars

OpenStudy (jdoe0001):

so if you want €2 , then 1.298 + 1.298 if you want €3, then 1.298 + 1.298 + 1.298 if you want €4, then 1.298 + 1.298 + 1.298 + 1.298 and so on

OpenStudy (jdoe0001):

now €2 , then 1.298 + 1.298 => 2 ( 1.298) €3, then 1.298 + 1.298 + 1.298 => 3 (1.298) €4, then 1.298 + 1.298 + 1.298 + 1.298 => 4 (1.298) and so on

OpenStudy (jdoe0001):

so what do you think it'll be for €75 in US dollars?

OpenStudy (anonymous):

do i divide?

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

well \(\begin{matrix} 2euros, \implies1.298 + 1.298 & \implies 2 ( 1.298)\\ 3euros, \implies 1.298 + 1.298 + 1.298 &\implies 3 (1.298)\\ 4euros, \implies1.298 + 1.298 + 1.298 + 1.298 & \implies 4 (1.298) \end{matrix}\)

OpenStudy (jdoe0001):

so what do you think it'll be for 75 Euros in US dollars?

OpenStudy (anonymous):

77.56?

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

@jim_thompson5910 help?

OpenStudy (jdoe0001):

\(\begin{matrix} \color{red}{2}euros, \implies1.298 + 1.298 & \implies \color{red}{2} \times ( 1.298)\\ \color{red}{3}euros, \implies 1.298 + 1.298 + 1.298 &\implies \color{red}{3} \times (1.298)\\ \color{red}{4}euros, \implies1.298 + 1.298 + 1.298 + 1.298 & \implies \color{red}{4} \times (1.298)\\ \color{red}{5}euros, \implies1.298 + 1.298 + 1.298 + 1.298+ 1.298 & \implies \color{red}{5} \times (1.298)\\ \color{red}{75}euros \implies 1.298 +.....+1.298& \implies \$\ \square? \end{matrix}\)

OpenStudy (jdoe0001):

is rather very straightforward

OpenStudy (anonymous):

oooooooooooooooooooooooooooohhhhhhhhhhhhhhhhhhhhhhhh thaaaaaaaaaaaaaaaaaaaaaaaaaannnnnnnnnnnnkkkkkkkkkkkkk yooooooooooooooooooooooooooou so much

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

97.35